OpenStudy (anonymous):

Which equation best represents the following scenario? Marino answered 3 more questions correctly on the quiz than Georgia. Marino answered ten questions correctly.

OpenStudy (anonymous):

@SiennaSurfer

OpenStudy (anonymous):

are there answers to choose from?

OpenStudy (anonymous):

3g =10 3 - g = 10 10 / g = 3 g + 3 = 10

OpenStudy (anonymous):

g + 3 = 10

OpenStudy (anonymous):

yep. its the last 1